OpenStudy (dreamsawakes):

The moon forms a right triangle with the Earth and the Sun during one of its phases, as shown below:

OpenStudy (kjkhjkh76876):

180(n−2)n=168

OpenStudy (dreamsawakes):

A scientist measures the angle x and the distance y between the Sun and the moon. Using complete sentences, explain how the scientist can use only these two measurements to calculate the distance between the Earth and the moon.

OpenStudy (dreamsawakes):

o-o wait what

OpenStudy (dreamsawakes):

Please help!

OpenStudy (hungergames13):

Designate the unknown distance between the earth and the sun as distance z. Therefore... tan x = y/z z(tan x) = y z = y/tan x

OpenStudy (dreamsawakes):

is that it?

OpenStudy (mathstudent55):

The tangent does not involve the hypotenuse.

OpenStudy (mathstudent55):

\(\sin \theta = \dfrac{opp}{hyp} \) \(\cos \theta = \dfrac{adj}{hyp} \) \(\tan \theta = \dfrac{opp}{adj} \)

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

We are looking for distance "a", the distance between the Earth and the Moon.

OpenStudy (mathstudent55):

In the triangle, side y is the hypotenuse. For angle x, side a is the opposite leg. The trig function that relates the opposite leg to the hypotenuse is the sine. \(\sin x = \dfrac{opp}{hyp} \) \(\sin x = \dfrac{a}{y}\) \(a = y \sin x\)

OpenStudy (dreamsawakes):

Thank you, I just got that.